# PipeOffload: Improving Scalability of Pipeline Parallelism with Memory Optimization

### v1

@@ 1 Introduction
Formally, for a single transformer layer, if we define $T_{o}$ as the round-trip time for its activation memory to be moved from device to host (D2H) and then host to device (H2D), $T_{c}$ as the time for its total forward and backward compute, then the ratio between them, denoted as $k=T_{o}/T_{c}$, is an important indicator on the proportion of the activation memory that can be offloaded without introducing significant efficiency drawbacks. Full activation memory offload is possible if $k\leq 1$. Without considering the trivial layers which can be recomputed with negligible overhead (e.g. Dropout, GeLU, LayerNorm), we can estimate the offload cost for the rest of the layers and subsequently estimate $k$. With sequence length ($s$), hidden size ($h$), PCI-E duplex bandwidth ($B_{o}$) and GPU compute bandwidth ($B_{c}$), $k$ is computed as: (Narayanan et al., 2021; Korthikanti et al., 2023)
@@
Note that the value of $k$ decreases as model size or sequence length increases. Figure 1 (left) demonstrates that the value of $k$ is surprisingly small under typical hidden dimension and sequence size settings, indicating that a significant portion of activation memory can be offloaded. An interesting observation is that when the hidden dimension exceeds 8k or the sequence length surpasses 16k, all activation memory can be offloaded with negligible overhead as shown in Figure 1 (right).
When $k>1$, offloading all activation memory is impossible without sacrificing throughput. Under this scenario, we resort to a partial offload where a subset of the activations gets offloaded. Similar to the rematerialization approach where trivial operations are preferred as they impose less computation burden, we perform selective offload that prioritize activations that yield the greatest reduction in peak memory usage. We introduce a general guideline for selective offload, always preferring activations with longer lifespan, namely longer gaps between forward and backward passes. Intuitively, the longer an activation remains in flight, the more it contributes to peak memory.
A widely adopted strategy to reduce pipeline bubbles involves placing multiple stages on the same device, as demonstrated in works such as ((Narayanan et al., 2021), (Qi et al., 2024), (Liu et al., 2023)). Notably, different stages has varying lifespan, potentially resulting in a more efficient better-than-linear reduction in peak memory usage. The efficiency of selective offload also hinges on the memory usage pattern of the pipeline schedule. For instance, when visualizing the memory usage patterns of interleaved 1F1B (Narayanan et al., 2021) and our PipeOffload method in Figure 2, it becomes apparent that offloading stage 0 in our method yields a 3/4 reduction in peak memory, whereas in the interleaved 1F1B, at most a 1/2 reduction can be achieved.
When memory offload is applied in practice, we need to further consider its interplay with other factors, especially the trade-off between memory and throughput. We extend the interleaving strategy into a generalized form, offering smooth memory reduction with minimal throughput loss. This approach provides flexibility to optimize performance based on specific system needs.
In the rest of this paper, we describe in detail the selective offload strategy in Section 2, we introduce a family of pipeline schedules that trade-off between throughput and peak memory in Section 3, we elaborate on implementation details in Section 4 and finally evaluate and compare the methods in Section 5.

### v2

@@ 1 Introduction
Formally, for a single transformer layer, if we define $T_{o}$ as the round-trip time for its activation memory to be moved from device to host (D2H) and then host to device (H2D), $T_{c}$ as the time for its total forward and backward compute, then the ratio between them, denoted as $k=T_{o}/T_{c}$, is an important indicator on the proportion of the activation memory that can be offloaded without introducing significant efficiency drawbacks. Full activation memory offload is possible if $k\leq 1$, as the offload operations can be fully overlapped by compute. Without considering the trivial layers which can be recomputed with negligible overhead (e.g. Dropout, GeLU, LayerNorm), we can estimate the offload cost for the rest of the layers and subsequently estimate $k$. With sequence length ($s$), hidden size ($h$), PCI-E duplex bandwidth ($B_{o}$) and GPU compute bandwidth ($B_{c}$), $k$ is computed as: (Narayanan et al., 2021; Korthikanti et al., 2023)
@@
Note that the value of $k$ decreases as model size or sequence length increases. Figure 1 (left) demonstrates that the value of $k$ is surprisingly small under typical hidden dimension and sequence size settings, indicating that a significant portion of activation memory can be offloaded. We observe that $k\leq 1$ when the hidden dimension exceeds 8k or the sequence length is greater than 16k; under these conditions, all activation memory can be offloaded with negligible overhead, as shown in Figure 1 (right).
When $k>1$, offloading all activation memory would lead to a decrease in throughput. In this scenario, we resort to a partial offload where a subset of the activations are offloaded. Similar to the rematerialization approach where trivial operations are preferred as they impose less computation burden, we perform selective offload that prioritize activations that yield the greatest reduction in peak memory usage. We introduce a general guideline for selective offload that always prefers activations with longer lifespan, i.e., longer gaps between forward and backward passes. Intuitively, the longer an activation remains in flight, the more it contributes to peak memory.
A widely adopted strategy to reduce pipeline bubbles involves placing multiple stages on the same device, as demonstrated in works such as (Narayanan et al., 2021), (Qi et al., 2024), and (Liu et al., 2023). Notably, different stages have varying lifespans, potentially resulting in a more efficient, better-than-linear reduction in peak memory usage. The efficiency of selective offload also depends on the memory usage pattern of the pipeline schedule. For instance, when visualizing the memory usage patterns of interleaved 1F1B (Narayanan et al., 2021) and our PipeOffload method in Figure 2, it is apparent that offloading stage 0 in our method yields a 3/4 reduction in peak memory, while in the interleaved 1F1B, at most a 1/2 reduction can be achieved.
When memory offload is applied in practice, we need to further consider its interplay with other factors, especially the trade-off between memory and throughput. We extend the interleaving strategy into a generalized form, offering smooth memory reduction with minimal throughput loss. This approach provides flexibility in optimizing performance based on specific system needs.
In the rest of this paper, we detail the selective offload strategy in Section 2, introduce a family of pipeline schedules that trade-off between throughput and peak memory in Section 3, elaborate on implementation details in Section 4, and finally evaluate and compare the methods in Section 5.
@@ 5.4 Comparing With Tensor Parallelism
5.5 Comparing With Other Pipeline Parallel Methods
We compare our approach with existing pipeline parallelism (PP) methods, particularly those aimed at reducing activation memory, including V-Min and V-Half from Qi et al. (2024), as well as activation re-materialization combined with 1F1B-I (shown as 1F1B-I-R). As shown in Figure 12, our methods achieve a more favorable Pareto frontier in the trade-off between memory usage and throughput. Through detailed analysis, we attribute the suboptimal performance of the V-Shaped schedule proposed in Qi et al. (2024) to significant disparities among $T_{F}$, $T_{B}$, and $T_{W}$ at large sequence lengths. These disparities violate the core assumption in Qi et al. (2024) that these timings are approximately equal, leading to degraded efficiency.
Figure 12: Pareto frontier compared with other methods.
5.6 Convergence Experiment
We compare the loss curves of our PO-H and PO-F implementations with 1F1B-I to validate correctness on a 5.4B model using 8 GPUs. As shown in Figure 13, the results exhibit nearly identical loss trajectories, providing strong evidence for convergence correctness. Notably, our methods are designed as exact algorithms with no compromises in accuracy.
Figure 13: Loss curves proves algorithmic correctness of our methods.
@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Large Language Model Training Systems. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.